# Approaching the thermodynamic limit of a bounded one-component plasma

### v1

@@ II.2 Equations of state and thermodynamic functions
In Sec. II.1, we have explored the ionic subsystem and derived the equation of state (EOS) (16) for it. Below, we will obtain EOS for the entire system, which is supposed to include a contribution from the background. Hence, this requires some physical idea of the background nature. We consider the simplest model of the background as a system of a large number $\bar{N}\to\infty$ of small charged pseudo-particles of the charge ${\bar{z}}e$ uniformly distributed in the volume $V_{R}$ such that $\bar{N}\bar{z}+Nz=0$ [50]. Here, the quantities related to the pseudo-particles are marked by bar. In this limit, it is obvious that both the pseudo-particle–pseudo-particle and pseudo-particle–ion coupling parameters, $\Gamma{(N/\bar{N})^{5/3}}$ and $\Gamma{(N/\bar{N})^{2/3}}$, respectively, vanish, and the background electrostatic energy is defined by Eq. (4). Now, we have to treat the system potential energy as a homogeneous function of the coordinates of the order 1-1 and write the equation of state for the entire system based on the virial theorem as
@@
where $p$ is the system pressure and $\bar{K}=(3/2)\bar{N}T$ is the pseudo-particle kinetic energy. As is well-known, the totally neutral system of classical Coulomb charges tends to collapse, and the main background function is to stabilize the entire system. This is reached by the ideal gas pressure of the pseudo-particles, which tends to infinity as $\bar{N}\to\infty$. However, one can renormalize this pressure by introducing the excess pressure $p\to p-2\bar{K}/3{V_{R}}$ [50], so that the compressibility factor is reduced to
@@
(see, cf. [40]). In terms of virials, Eq. (20) with $\bar{K}=0$ can be written as
$p=\displaystyle{N\over{{V_{R}}}}\left({T+\displaystyle{1\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{i}}}\right\rangle}+\displaystyle{1\over{3N}}\sum\limits_{j=1}^{\bar{N}}{\left\langle{{{{\bf{\bar{r}}}}_{j}}{{{\bf{\bar{f}}}}_{j}}}\right\rangle}}\right),$ (22)
@@
$p_{i}=\displaystyle{N\over{{V_{R}}}}\left({T+\displaystyle{1\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{i}}}\right\rangle}}\right)$ (23)
@@
${Z_{i}}=1+\displaystyle{\Gamma\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{i}}}\right\rangle}.$ (24)
The pressures $p$ and $p_{i}$ and corresponding compressibility factors $Z$ and $Z_{i}$ differ by the third term in parenthesis on the right-hand side of (22), which cannot be determined in any simulation. Therefore, although the pressure (22) must be fully compatible with its thermodynamic definition because it leads to (21), it is impossible to obtain it from simulations. Incompatibility of the pressure defined thermodynamically with that obtained from the virial was noted in study [63], where a different interpretation for such discrepancy was formulated. Note that the ionic pressure (15) and (23) are compatible with thermodynamics and can be determined directly from MD simulation. In addition, it is the only pressure that positively has physical meaning.
If we treat the ions in the field of background as a separate system, we can define the excess internal energy of the ions $\varepsilon_{i\mathrm{ex}}(\Gamma,N)$ per ion. Since at $T=0$, $u_{i}=u_{p}+u_{b}=u_{0}-u_{c}$, relative to this reference state,
$\varepsilon_{i\mathrm{ex}}(\Gamma,N)=u(\Gamma,N)-u_{0}$ (25)
depends on the total system electrostatic energy. Note that $\varepsilon_{i\rm{ex}}(\Gamma,N)$ has a thermodynamic limit, while $u_{p}+u_{b}$ does not. In the thermodynamic limit, at $\Gamma>{\Gamma_{\mathrm{m}}}$, where $\Gamma={\Gamma_{\mathrm{m}}}$ is the melting point of OCP, the harmonic lattice model is applicable [66, 26], in which
$\varepsilon_{i\rm{ex}}(\Gamma,\infty)=\displaystyle{3\over 2\Gamma}+\displaystyle{\beta_{2}\over\Gamma^{2}},$ (26)
with the coefficient $\beta_{2}$ defining anharmonicity. The thermodynamic limit of the excess ionic Helmholtz free energy per ion that can be calculated from (25) using (see, e.g. [35]),
${f_{i\mathrm{ex}}}(\Gamma)=-\int\limits_{\Gamma}^{\infty}{\displaystyle{{{\varepsilon_{i\mathrm{ex}}}(\Gamma^{\prime},\infty)}\over{\Gamma^{\prime}}}}\,d\Gamma^{\prime},$ (27)
is a function of $u_{\infty}(\Gamma)$. If $\Gamma_{\rm{m}}$ and $\beta_{2}$ are known, the excess ionic Helmholtz free energy, which coincides with this quantity for the entire system, can be derived using (26):
${f_{i\mathrm{ex}}}({\Gamma_{\mathrm{m}}})=-\displaystyle{3\over{2{\Gamma_{\mathrm{m}}}}}-\displaystyle{\beta_{2}\over{2\Gamma_{\mathrm{m}}^{2}}}.$ (28)
@@ III.1 Simulation method
An advantage of the BOCP simulation is absence of diverging quantities due to finiteness of $N$, which makes it possible to determine all quantities of interest including $u$, $u_{p}$, and $u_{b}$. In contrast, calculation of the total potential energy $u$ for the OCP with PBC requires application of special summation methods such as the Ewald potential [56, 40], angular-averaged Ewald potential [47], particle-particle particle-mesh method (PPPM) [67], etc., while for BOCP, this quantity can be determined ab initio from direct summation of the interparticle Coulomb energies and the particle–background energy. Moreover, since no summation method for $u_{p\mathrm{ex}}$ and $u_{p\mathrm{ex}}$ for OCP has been developed, BOCP simulation remains a unique method for estimation of these quantities. Obvious shortcomings of the BOCP simulation are enormous computational efforts when no acceleration procedure is possible and the necessity to simulate a whole spectrum of the BOCP sizes to enable the extrapolation to the thermodynamic limit.
MD simulation of ions motion within the background sphere is performed. Introduction of the dimensionless dynamic quantities by replacement for the time $t$ $\to t/{\tau_{0}}$, the radius-vector of the $i$th ion ${{\bf{r}}_{i}}\to{{\bf{r}}_{i}}/{r_{s}}$, and the force acting on this ion ${{\bf{f}}_{i}}\to({r_{s}}/\varepsilon){{\bf{f}}_{i}}=\left({\tau_{0}^{2}/m{r_{s}}}\right){{\bf{f}}_{i}}$ makes it possible to write the equation of the ion motion in the form [62]
${{\bf{\ddot{r}}}_{i}}={{\bf{f}}_{zi}}+{{\bf{f}}_{bi}}+{{\bf{f}}_{Li}},$ (29)
@@
${{\bf{f}}_{zi}}=\sum\limits_{j\neq i}{\displaystyle{{{\bf{r}}_{i}}-{{\bf{r}}_{j}}\over r_{ij}^{3}}}$ (30)
@@
${\bf f}_{bi}=-{\bf r}_{i}$ (31)
@@
${{\bf{f}}_{Li}}=-\gamma{{\bf{\dot{r}}}_{i}}+{{\bf{f}}_{\mathrm{st}i}}$ (32)
is the force acting from the Langevin thermostat. Here, $\gamma$ is the particle friction coefficient, ${\bf{f}}_{\mathrm{st}i}$} is a random force with the Gaussian distribution such that
@@
where $\tau_{\mathrm{st}}$} is the autocorrelations decay time, which was set equal to the time integration step $\tau$, $\left\langle{{v^{2}}}\right\rangle\equiv\left\langle{v_{i}^{2}}\right\rangle$ is the mean squared ion velocity, which is independent of $i$ at equilibrium, and we take into account that in equilibrium, $\left\langle{{v^{2}}}\right\rangle=3T=3/\Gamma$.
@@
A contact with the thermostat in the course of simulation reduces sharply the system equilibration time and ensures a very small deviation of $\left\langle{{v^{2}}}\right\rangle/3$ from the desired value $T$ thus decreasing sharply the computation time. However, the question arises whether the inclusion of $f_{Li}$} in the total force can distort the virial. Below, we demonstrate that no distortion can be observed because $\sum\nolimits_{i=1}^{N}{{{\bf{r}}_{i}}{{\bf{f}}_{Li}}}=0$. First, consider the case of ideal gas ($\Gamma\to 0$), then Eq. (29) can be rewritten as
${{\bf{\ddot{r}}}_{i}}=-\gamma{{\bf{\dot{r}}}_{i}}+{{\bf{f}}_{\mathrm{st}i}}.$ (34)
@@
We take the dot products of both sides of (34) and ${\bf{r}}_{i}$, average this equation, and take into account that $\left\langle{{{\bf{r}}_{i}}{{{\bf{\ddot{r}}}}_{i}}}\right\rangle=d\left\langle{{{\bf{r}}_{i}}{{{\bf{\dot{r}}}}_{i}}}\right\rangle/dt-\left\langle{{v^{2}}}\right\rangle$ and
${d\over{dt}}\left\langle{{{\bf{r}}_{i}}{{{\bf{\dot{r}}}}_{i}}}\right\rangle=\displaystyle{1\over 2}\displaystyle{{{d^{2}}}\over{d{t^{2}}}}\left\langle{r_{i}^{2}}\right\rangle=\displaystyle{D\over 2}\displaystyle{{{d^{2}}t}\over{d{t^{2}}}}=0,$ (35)
@@
where we imply that an ion is involved in the Brownian motion when $\left\langle{r_{i}^{2}}\right\rangle=Dt$ with the diffusion coefficient $D={\tau_{\mathrm{st}}}\left\langle{{v^{2}}}\right\rangle$. Bearing this in mind we infer from Eqs. (33) and (34) that $\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{\mathrm{st}i}}}\right\rangle=0$, so that the virial
@@
$\sum\limits_{k=1}^{\infty}{\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{Li}}}\right\rangle}=-\gamma\sum\limits_{k=1}^{\infty}{\left\langle{{{\bf{r}}_{i}}{{{\bf{\dot{r}}}}_{i}}}\right\rangle}+\sum\limits_{k=1}^{\infty}{\left\langle{{{\bf{r}}_{i}}{{\bf{f}}_{\mathrm{st}}}}\right\rangle}=0.$ (36)
vanishes exactly.
@@
Next, consider the case of a strongly coupled system, $\Gamma\to\infty$, for which we use again the ion sphere model (Sec. II.1). Now the ion equation of motion includes the electrostatic force ${\bf{f}}=-{\bf{r}}$: ${\bf{\ddot{r}}}=-{\bf{r}}-\gamma{\bf{\dot{r}}}+{{\bf{f}}_{\mathrm{st}}}$, where we omit subscripts. In contrast to the case of ideal gas treated above, an ion in a strongly coupled system is involved in the Ornstein–Uhlenbeck process in its microscopic sphere, so that $\left\langle{{\bf{r}}{\bf{\dot{r}}}}\right\rangle=(1/2)d\left\langle{{r^{2}}}\right\rangle/dt=0$, and therefore, $\left\langle{{\bf{r}}{\bf{\ddot{r}}}}\right\rangle=-\left\langle{{v^{2}}}\right\rangle$. We calculate the dot product of both sides of the ion equation of motion and $\bf{r}$}. With due regard for the equality $\left\langle{{v^{2}}}\right\rangle=\left\langle{{r^{2}}}\right\rangle$ valid at equilibrium, we derive the same result as above: ${\bf{r}}{{\bf{f}}_{L}}=0$ [62].
@@
In the intermediate case of moderate $\Gamma$, the ion is involved in both the Brownian and Ornstein–Uhlenbeck process, i.e., its motion is characterized by two time scales. Such process is quite similar to that observed in study [68]. Test runs showed that irrespective of $\Gamma$, all results were absolutely insensitive to the choice of $\gamma$, which confirms the conclusions made above (note that it must be true solely at equilibrium). For all BOCP simulations, we borrowed $\gamma\simeq 0.09$ from earlier studies [20, 64, 62].
@@
We apply the specular reflection boundary conditions for the equation of motion (29), i.e., when an ion crosses the surface of a sphere of the radius $R_{b}$ then its velocity ${\bf{v}}_{i}$ is transformed as
${{\bf{v}}^{\prime}_{i}}={{\bf{v}}_{i}}-\displaystyle{{2{{\bf{r}}_{i}}{{\bf{v}}_{i}}}\over{r_{i}^{2}}}{{\bf{r}}_{i}}.$ (37)
@@
Due to finiteness of the time integration step, the ion velocity can be transformed only at the point where $r_{i}>R_{b}$, i.e., the ion density distribution $f_{c}(r)$ is never sigmoid-shaped but its vanishing at $r\gtrsim R_{b}$ is smooth. Hence, if we would set $R_{b}=R$ then some artificial charged layer was formed beyond the surface of compensating background, which leads to violation of the charge neutrality inside BOCP. This effect cannot be eliminated completely but it can be minimized if $R_{b}$ coincides with the ion equimolar radius. Since the average distance at which an ion propagates during the time $\tau$ is $\tau/\sqrt{2\pi\Gamma}$, the shift $R_{b}=R-\tau/\sqrt{8\pi\Gamma}$ minimizes the boundary effect. Note that variation of such shift by several times does not change the results principally but it increases significantly the smallest size beginning from which the size dependence $u(\Gamma)$ conforms to Eq. (19) thus indicating the increase of the surface effect. Additionally, at each time step, the ion velocity was normalized to $\sqrt{2/\Gamma}$ if it exceeded $\sqrt{25/\Gamma}$.
@@
$\Gamma$} $\tau$} Range of NN Range of ${10^{4}}N_{\tau}$
@@
The choice of time integration step depends on $\Gamma$, and $\tau(\Gamma)$ is an increasing function due to slow-down of the thermal motion. If $\tau$ is too small then too large number of time steps $N_{\tau}$ that exceeds the computational resources is required to equilibrate the system. For too large $\tau$, the accuracy of the used numerical integration scheme becomes insufficient, which results in noticeable errors in determined parameters. Thus, an optimum $\tau$ was determined for each $\Gamma$ from the condition that the determined parameters are almost insensitive to variation of $\tau$ by several times (Table 1).
MD simulation of BOCP was performed for the set of BOCP sizes $N=2500$, 3540, 5000, 7070, 10000, 14000, 21200, 30000, and 50000 at different $\Gamma$. Due to limitation of computational resources, the maximum $N$ is defined by the minimum $N_{\tau}$, for which variation of the determined quantity is much less than its estimated error. Hence, in some cases, simulations were performed in a limited range of $N$. The ranges of $N$ and $N_{\tau}$ estimated in such a way for different $\Gamma$’s are also listed in Table 1. For the sizes within the range limits, $N_{\tau}\propto N^{-2}$. The upper bound of $N$ at $\Gamma=200$ and 210 is a consequence of higher fluctuations in the system, where a transitional regime from fluid to solid (fluid core and solid shells) takes place. At $\Gamma=650$ and 1000, the BOCP equilibration time is sharply increased due to the decrease in the ion self-diffusion coefficient for a solid system. For $\Gamma=0.03$, the lower bound of $N$ is due to the fact that the size dependence $u(\Gamma)$ does not conform to Eq. (19) at $N<3540$.
First, the positions of $N$ ions were initialized by the homogeneous random spatial distribution in a sphere of the radius $R_{b}$. Then Eqs. (29) were integrated numerically using the Verlet algorithm. After the relaxation of the system to equilibrium state after $N_{\tau}/2$ time steps, data were collected every 100 time steps. For the calculation of $f_{p}(r)$ (7), an ion closest to the BOCP center was selected with the same time interval. The data for different quantities were averaged, and the error was estimated specifically for each quantity (Sec. III.2). Eventually, obtained data were used for determination of the OCP properties in the thermodynamic limit.

### v2

@@ II.2 Equations of state and thermodynamic functions
In Sec. II.1, we have explored the ionic subsystem and derived EOS (16) for it. Below, we will obtain the EOS for the entire BOCP, which is supposed to include a contribution from the background, and demonstrate that the pressure for the entire BOCP is principally different from the ionic pressure. This requires some physical model of the background. We consider the simplest model of the background as a system of a large number $\bar{N}\to\infty$ of small charged pseudo-particles of the charge ${\bar{z}}e$ uniformly distributed in the volume $V_{R}$ such that $\bar{N}\bar{z}+Nz=0$ [61]. Here, the quantities related to the pseudo-particles are marked with bars. In this limit, it is obvious that both the pseudo-particle–pseudo-particle and pseudo-particle–ion coupling parameters, $\Gamma{(N/\bar{N})^{5/3}}$ and $\Gamma{(N/\bar{N})^{2/3}}$, respectively, vanish, and the background electrostatic energy is defined by Eq. (4). Now, we have to treat the system potential energy as a homogeneous function of the order 1-1 in the coordinates and write the EOS for entire system based on the virial theorem as
@@
where $p$ is the system pressure and $\bar{K}=(3/2)\bar{N}T$ is the pseudo-particle kinetic energy. As is well known, the totally neutral system of classical Coulomb charges tends to collapse, and the main background function is to stabilize the entire system. This is reached by the ideal gas pressure of the pseudo-particles, which tends to infinity as $\bar{N}\to\infty$. However, one can renormalize this pressure by introducing the excess pressure $p\to p-2\bar{K}/3{V_{R}}$ [61], so that the compressibility factor is reduced to
@@
(cf. [51]). In terms of virials, Eq. (20) with $\bar{K}=0$ can be written as
$p=\displaystyle{N\over{{V_{R}}}}\left({T+\displaystyle{1\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{i}}}\right\rangle}+\displaystyle{1\over{3N}}\sum\limits_{j=1}^{\bar{N}}{\left\langle{{{{\mathbf{\bar{r}}}}_{j}}{{{\mathbf{\bar{f}}}}_{j}}}\right\rangle}}\right),$ (22)
@@
$p_{i}=\displaystyle{N\over{{V_{R}}}}\left({T+\displaystyle{1\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{i}}}\right\rangle}}\right)$ (23)
@@
${Z_{i}}=1+\displaystyle{\Gamma\over{3N}}\sum\limits_{i=1}^{N}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{i}}}\right\rangle}.$ (24)
The pressures $p$ and $p_{i}$ and corresponding compressibility factors $Z$ and $Z_{i}$ differ by the third term in parentheses on the right-hand side of (22), which cannot be determined in any simulation. Therefore, although the pressure (22) must be fully compatible with its thermodynamic definition because it leads to (21), it cannot be obtained from simulations. The resulting incompatibility of entirely different quantities, $p$ and $p_{i}$ is worth mentioning. It is $p_{i}$ that is actually determined in both MD and MC simulations, therefore, $p_{i}$ cannot be compared with the thermodynamically defined system pressure $p$ regardless of whether it is renormalized, despite the fact that both the thermodynamic approach and virial theorem are suitable for the treated system.
If we treat the ions in the field of background as a separate system, we can define the excess internal energy of ions $\varepsilon_{\mathrm{ex}}(\Gamma,N)$ per ion, which coincides with the excess internal energy of the entire system (Sec. II.1). If we choose $\Gamma=\infty$ as a reference state, then
$\varepsilon_{\mathrm{ex}}(\Gamma,N)=u(\Gamma,N)-u(\infty,N).$ (25)
In the thermodynamic limit, $\varepsilon_{\mathrm{ex}}(\Gamma)=u_{\infty}(\Gamma)-u_{0}$. At $\Gamma>{\Gamma_{\mathrm{m}}}$, where $\Gamma={\Gamma_{\mathrm{m}}}$ is the melting point of OCP, the harmonic lattice model supplemented by anharmonicity correction is applicable [77, 34], which yields
$\varepsilon_{\rm{ex}}(\Gamma)=\displaystyle{3\over 2\Gamma}+\displaystyle{\beta_{2}\over\Gamma^{2}}$ (26)
with the coefficient $\beta_{2}$ defining anharmonicity. The thermodynamic limit of the excess Helmholtz free energy per ion of the entire system, which is equal to that of the ionic subsystem, can be derived from $\varepsilon_{\rm{ex}}(\Gamma)$ (i.e., from $u_{\infty}(\Gamma)$, see, e.g., [46]):
${f_{\mathrm{ex}}}(\Gamma)=-\int\limits_{\Gamma}^{\infty}{\displaystyle{{{\varepsilon_{\mathrm{ex}}}(\Gamma^{\prime})}\over{\Gamma^{\prime}}}}\,d\Gamma^{\prime}.$ (27)
Then at the melting point, we obtain from (26) and (27)
${f_{\mathrm{ex}}}({\Gamma_{\mathrm{m}}})=-\displaystyle{3\over{2{\Gamma_{\mathrm{m}}}}}-\displaystyle{\beta_{2}\over{2\Gamma_{\mathrm{m}}^{2}}}.$ (28)
@@ III.1 Simulation method
An advantage of the BOCP simulation is the absence of diverging quantities due to finiteness of NN, which makes it possible to determine all quantities of interest including $u$, $u_{p}$, and $u_{b}$. In contrast, calculation of the total potential energy $u$ for the OCP with PBC requires application of special summation methods such as the Ewald potential [68, 51], angular-averaged Ewald potential [58], particle-particle particle-mesh method (PPPM) [78], etc., while for BOCP, this quantity can be determined ab initio from direct summation of the interparticle Coulomb energies and the particle–background energy. Moreover, since no summation method for $u_{p\mathrm{ex}}$ and $u_{b\mathrm{ex}}$ for the OCP has been developed, BOCP simulation remains a unique method for the estimation of these quantities. Obvious shortcomings of the BOCP simulation are enormous computational efforts when no acceleration procedure is possible and the necessity to simulate a whole spectrum of the BOCP sizes to enable the extrapolation to the thermodynamic limit.
MD simulation of ions motion within the background sphere is performed. Introduction of the dimensionless dynamic quantities by replacing the time $t$ $\to t/{\tau_{0}}$, the radius-vector of the $i$th ion ${{\mathbf{r}}_{i}}\to{{\mathbf{r}}_{i}}/{r_{s}}$, and the force acting on this ion ${{\mathbf{f}}_{i}}\to({r_{s}}/\varepsilon){{\mathbf{f}}_{i}}=\left({\tau_{0}^{2}/m{r_{s}}}\right){{\mathbf{f}}_{i}}$ makes it possible to write the equation of the ion motion in the form [73]
${{\mathbf{\ddot{r}}}_{i}}={{\mathbf{f}}_{zi}}+{{\mathbf{f}}_{bi}}+{{\mathbf{f}}_{Li}},$ (29)
@@
${{\mathbf{f}}_{zi}}=\sum\limits_{j\neq i}{\displaystyle{{{\mathbf{r}}_{i}}-{{\mathbf{r}}_{j}}\over r_{ij}^{3}}}$ (30)
@@
${{\mathbf{f}}_{bi}}=\left\{{\begin{array}[]{cc}{-{{\mathbf{r}}_{i}},}&{\mbox{ }{r_{i}}\leq R,}\\ {-N\displaystyle{{{{\mathbf{r}}_{i}}}\over{r_{i}^{3}}},}&{\mbox{ }{r_{i}}>R}\end{array}}\right$. (31)
@@
${{\mathbf{f}}_{Li}}=-\gamma{{\mathbf{\dot{r}}}_{i}}+{{\mathbf{f}}_{\mathrm{st}i}}$ (32)
is the force acting from the Langevin thermostat. Here, $\gamma$ is the particle friction coefficient, ${\mathbf{f}}_{\mathrm{st}i}$} is a random force with the Gaussian distribution such that
@@
where $\tau_{\mathrm{st}}$} is the autocorrelation decay time, which was set equal to the time integration step $\tau$, $\left\langle{{v^{2}}}\right\rangle\equiv\left\langle{v_{i}^{2}}\right\rangle$ is the mean squared ion velocity, which is independent of $i$ at equilibrium, and we take into account that in equilibrium, $\left\langle{{v^{2}}}\right\rangle=3T=3/\Gamma$.
@@
Contact with the thermostat in the course of simulation reduces sharply the system equilibration time and ensures a very small deviation of $\left\langle{{v^{2}}}\right\rangle/3$ from the desired $T$ thus decreasing substantially the computation time. However, the question arises whether the inclusion of ${\mathbf{f}}_{Li}$ in the total force can distort the virial. Below, we demonstrate that no distortion can be observed because $\sum\nolimits_{i=1}^{N}{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{Li}}}=0$. First, consider the case of ideal gas ($\Gamma\to 0$), in which case Eq. (29) can be rewritten as
${{\mathbf{\ddot{r}}}_{i}}=-\gamma{{\mathbf{\dot{r}}}_{i}}+{{\mathbf{f}}_{\mathrm{st}i}}.$ (34)
@@
We multiply both sides of (34) by ${\mathbf{r}}_{i}$, average this equation, and take into account that $\left\langle{{{\mathbf{r}}_{i}}{{{\mathbf{\ddot{r}}}}_{i}}}\right\rangle=d\left\langle{{{\mathbf{r}}_{i}}{{{\mathbf{\dot{r}}}}_{i}}}\right\rangle/dt-\left\langle{{v^{2}}}\right\rangle$ and
${d\over{dt}}\left\langle{{{\mathbf{r}}_{i}}{{{\mathbf{\dot{r}}}}_{i}}}\right\rangle=\displaystyle{1\over 2}\displaystyle{{{d^{2}}}\over{d{t^{2}}}}\left\langle{r_{i}^{2}}\right\rangle=\displaystyle{D\over 2}\displaystyle{{{d^{2}}t}\over{d{t^{2}}}}=0,$ (35)
@@
where we imply that an ion is involved in the Brownian motion when $\left\langle{r_{i}^{2}}\right\rangle={\mathcal{D}}t$, where, as it follows from (33), ${\mathcal{D}}={\tau_{\mathrm{st}}}\left\langle{{v^{2}}}\right\rangle$. Bearing this in mind we infer from Eqs. (33) and (34) $\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{\mathrm{st}i}}}\right\rangle=0$, so that the virial vanishes:
@@
$\sum\limits_{i=1}^{\infty}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{Li}}}\right\rangle}=-\gamma\sum\limits_{i=1}^{\infty}{\left\langle{{{\mathbf{r}}_{i}}{{{\mathbf{\dot{r}}}}_{i}}}\right\rangle}+\sum\limits_{i=1}^{\infty}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{\mathrm{st}}}}\right\rangle}=0.$ (36)
@@
Next, consider the case of a strongly coupled system, $\Gamma\to\infty$, for which we use again the ion sphere model (Sec. II.1). Now the ion equation of motion includes the electrostatic force ${\mathbf{f}}=-{\mathbf{r}}$: ${\mathbf{\ddot{r}}}=-{\mathbf{r}}-\gamma{\mathbf{\dot{r}}}+{{\mathbf{f}}_{\mathrm{st}}}$, where we omit subscripts. In contrast to the case of ideal gas treated above, an ion in a strongly coupled system is involved in the Ornstein–Uhlenbeck process in its microscopic sphere, so that $\left\langle{{\mathbf{r}}{\mathbf{\dot{r}}}}\right\rangle=(1/2)d\left\langle{{r^{2}}}\right\rangle/dt=0$, and therefore, $\left\langle{{\mathbf{r}}{\mathbf{\ddot{r}}}}\right\rangle=-\left\langle{{v^{2}}}\right\rangle$. We multiply both sides of the ion equation of motion by $\mathbf{r}$} with due regard for the equality $\left\langle{{v^{2}}}\right\rangle=\left\langle{{r^{2}}}\right\rangle$ valid at equilibrium to derive the same result as above: ${\mathbf{r}}{{\mathbf{f}}_{L}}=0$ [73].
@@
In the intermediate case of moderate $\Gamma$, the ion is involved in both the Brownian and Ornstein–Uhlenbeck processes, i.e., its motion is characterized by two time scales. Such process is quite similar to that observed in study [79]. Test runs showed that irrespective of $\Gamma$, all results were absolutely insensitive to the choice of $\gamma$, which confirms the conclusions made above (note that this must be true solely at equilibrium). For all BOCP simulations, we take $\gamma\simeq 0.09$ from earlier studies [28, 75, 73].
@@
We apply the specular reflection boundary conditions for the equation of motion (29), i.e., when an ion crosses the surface of a sphere of the radius $R_{b}$ then its velocity ${\mathbf{v}}_{i}$ is transformed as
${{\mathbf{v}}^{\prime}_{i}}={{\mathbf{v}}_{i}}-\displaystyle{{2{{\mathbf{r}}_{i}}{{\mathbf{v}}_{i}}}\over{r_{i}^{2}}}{{\mathbf{r}}_{i}}.$ (37)
@@
Due to the finiteness of the time integration step, the ion velocity can be transformed only at the point where $r_{i}>R_{b}$, i.e., the ion density distribution $f_{c}(r)$ is never step-like but its vanishing at $r>R_{b}$ is smooth. Were we set $R_{b}=R$ then some artificial charged layer would form beyond the surface of compensating background, which violates the charge neutrality inside the BOCP. This effect cannot be eliminated completely but it can be minimized by slightly reducing $R_{b}$ so that it coincides with the ion equimolar radius. Since the average distance at which an ion propagates during the time step $\tau$ is $\tau/\sqrt{2\pi\Gamma}$, the shift $R_{b}=R-\tau/\sqrt{8\pi\Gamma}$ minimizes the boundary effect. Note that increasing or decreasing the shift by several times does not significantly change the results but the BOCP size beginning from which the size dependence $u(\Gamma)$ corresponds to Eq. (19) increases noticeably thus indicating the increase of the boundary effect. Additionally, at each time step and for each ion, the components of ion velocity were scaled such that its modulus was set to $\sqrt{2/\Gamma}$ if it exceeded $\sqrt{25/\Gamma}$.
@@
$\Gamma$} $\tau$} Range of NN Range of $N_{\tau}\times 10^{-4}$
@@
The choice of time integration step depends on $\Gamma$, and $\tau(\Gamma)$ is an increasing function due to slowdown of the thermal motion. If $\tau$ is too small then too large number of time steps $N_{\tau}$ that exceeds the computational resources is required to equilibrate the system. For too large $\tau$, the accuracy of the used numerical integration scheme becomes insufficient, which leads to noticeable errors in the parameters of interest. Thus, an optimum $\tau$ was determined for each $\Gamma$ from the condition that the determined parameters are almost insensitive to variation of $\tau$ by several times (Table 1).
MD simulation of BOCP was performed for the set of BOCP sizes $N=2500$, 3540, 5000, $70$, 10000, 14000, 21200, 30000, and 50000 at different $\Gamma$. Due to the limitation of computational resources, the maximum NN is defined by the minimum $N_{\tau}$, for which variation of the determined quantity is much less than its estimated error. Thus, in some cases, simulations were performed in a limited range of NN. The ranges of NN and $N_{\tau}$ estimated in such a way for different $\Gamma$’s are also listed in Table 1. For the sizes within the range limits, $N_{\tau}\propto N^{-2}$. The upper bound of NN at $\Gamma=200$ and 210 is a consequence of higher fluctuations in the system, where a transitional regime from fluid to solid (fluid core and solid shells) takes place. At $\Gamma=650$ and 1000, the BOCP equilibration time increases sharply due to the decrease in the ion self-diffusion coefficient for a solid system. For $\Gamma=0.03$, the lower bound of NN is due to the fact that the size dependence $u(\Gamma)$ does not correspond to Eq. (19) at $N<3540$.
First, the positions of NN ions were initialized using the homogeneous random spatial distribution in a sphere of the radius $R_{b}$. Then Eqs. (29) were integrated numerically using the Verlet algorithm. After the relaxation of the system to equilibrium over $N_{\tau}/2$ time steps, data were collected every 100 time steps. For the calculation of $f_{p}(r)$ (7), an ion closest to the BOCP center was selected at the same time interval. The data for different quantities were averaged, and the error was estimated specifically for each quantity (Sec. III.2). Eventually, the obtained data were used for the determination of the OCP properties in the thermodynamic limit.